OpenStudy (anonymous):

Find the indicated probability. If P(B) = 0.5, P(A or B) = 0.3, and P(A and B) = 0.6. P(A) =

OpenStudy (anonymous):

P(A) = x P(A or B) = P(A) + P(B) - P(A and B) 0.3 = x + 0.5 - 0.6 0.3 = x - 0.1 0.3 + 0.1 = x 0.4 = x x = 0.4 So P(A) = 0.4

OpenStudy (anonymous):

hope this help :)